OpenStudy (anonymous):

Yara rolls a standard die and flips a coin. What is the probability that she gets an even number and heads.

OpenStudy (anonymous):

i forgot to add the heads part

OpenStudy (mani_jha):

The total no of possible cases are 12. \[(1, head),(1,tail),(2,head)...........(6, head), (6, tail)\] The no of even number and heads is 3 \[(2, head), (4,head), (6, head)\] So it is 1/4

OpenStudy (anonymous):

\[\frac{1}{2}\times \frac{1}{2}=\frac{1}{4}\]